# Single-shot volumetric fluorescence imaging with neural fields

### v2

@@ 3.1 Simulation evaluation of system performance
We first assessed the resolution of the QuadraPol PSF by simulating images of two closely positioned point sources. To determine the resolution, we applyed the Rayleigh criterion, which requires at least a 20% dip between the peaks of the two reconstructed spots. We used a high spatial sampling rate to evaluate the resolution more accurately. In the lateral direction, the QuadraPol PSF achieves an in-focus resolution of 4.3 µm. However, this resolution rapidly degrades to approximately 7 µm with a ±$\pm 0.3$ mm defocus. The resolution remains under 9 µm with a defocus extending up to 1.8 mm (Figure 4(a)). To optimally adapt to these resolutions in our experimental setup, we used a camera sampling rate that corresponds to a resolution limit of 7.4 µm, achieving a DOF of $\sim$4 mm with relatively uniform lateral resolution. This extended DOF is notably longer than the 0.26 mm DOF achievable with a standard PSF.
@@
Additionally, we compared our neural fields reconstruction algorithm with the RL deconvolution. For a fair comparison between our neural fields method and the RL deconvolution, both methods were provided with the same theoretical PSFs, i.e., the optical system is ideal and free from aberrations. Thus, the initialization of the neural fields method (Figure 3(a)) employs theoretical PSF across the entire image volume in the simulations. While both algorithms perform similarly under ideal conditions, neural fields exhibit significantly improved robustness with noisy images (Figure 4(a) and Supplementary Figure S2). For instance, at a low signal level where a total of $3.16\times 10^{4}$ photons are captured from both emitters, the RL deconvolution results in a highly noisy reconstruction, whereas the neural fields can still distinctly resolve the two spots and offers a two-fold resolution improvement.
@@
The axial resolution within the 4 mm DOF ranges from 230 to 250 µm, except for positions close to the focal plane ($|z|<0.1$ mm), where it degrades to 405 µm (Figure 4(b)). This degradation is attributed to the symmetry which causes gradual depth gradient of the in-focus QuadraPol PSF. However, this symmetry is disrupted in the presence of aberrations; when aberrated PSFs were simulated, an improvement in axial resolution near the focal plane was observed (Supplementary Figure S3). We also experimentally demonstrated this improvement in Section 3.2. Similar to the case of lateral direction, using neural fields leads to a better axial resolution compared to RL deconvolution under noisy conditions, although the improvement is less pronounced than in the lateral direction. Again, we emphasize that this comparison is under ideal conditions and does not account for forward model mismatches that may occur in experimental settings. The space-bandwidth product (SBP) of our system is $\sim$5.2 million voxels over the 4 mm depth range. Note that the SBP for the biological experiments in later sections is greater than the value reported here. We are not confined to the 4 mm depth range, as we can still reconstruct the object with slightly reduced resolution at greater defocus distances.
Figure 4: Performance evaluation of the QuadraPol PSF using simulated data. (a) Lateral and (b) axial resolutions determined by the Rayleigh criterion as functions of axial position and signal level. Images show representative noisy data and reconstruction cross-sections using RL deconvolution and neural fields. Shaded areas represents the diffraction ($\lambda/2\mathrm{NA}$) and sampling limits. Scale bar: 20 µm in $xy$ view and 200 µm in $xz$ view. (c-e) Reconstruction results for various line structures; insets show simulated images. Parameters for simulating the QuadraPol PSF ($\mathrm{NA}=0.52$, magnification $M=-5.2$, camera pixel size $2.2$ µm) are adjusted to match those reported for Miniscope3D [11]. Scale bar: 100 µm; color bar: height in µm. (f,g) Object and images using the polarized spiral (PS2F) and QuadraPol PSFs with (f) vertical and (g) horizontal lines at $z=\pm 0.35$ mm. Scale bar: 50 µm; color bar: AoLP. Insets show the PSFs. (h) Comparison between the QuadraPol PSF and other SVF techniques.
@@
One key advantage of the QuadraPol PSF is its ability to resolve denser objects more effectively. By separating polarization channels, the compact QuadraPol PSF reduces the mixing of information from different 3D positions, thus relaxing the sparsity constraint in the reconstruction algorithm compared to other SVF techniques such as the Miniscope3D [11]. To demonstrate this, we simulated three different line structures and generated images using both the Miniscope3D PSF and the QuadraPol PSF, analyzing them with the same RL deconvolution algorithm without sparsity constraints. For a small structure where the raw image using the Miniscope3D shows minimal overlap, both methods accurately resolve the structure (Figure 4(c)). Additionally, the reconstruction quality of the Miniscope3D for defocused axial positions is superior due to its extended DOF. However, for larger objects where there is substantial information mixture in the raw Miniscope3D images (Figure 4(d,e)), it struggles to accurately recover the structure. In contrast, the QuadraPol PSF still resolves the object effectively with significantly fewer artifacts. This enhanced performance is similarly observed with other dense objects (Supplementary Figure S4).
@@
A comparison of the three aspects mentioned above between the QuadraPol PSF and other SVF techniques is shown in Figure 4(h). The SBP of the QuadraPol PSF surpasses that of the Fourier lightfield microscope, as it does not compromise the field of view (FOV). However, despite the resolution improvements achieved using neural fields, its SBP does not match that of the MiniScope3D or the Fourier DiffuserScope, primarily due to the QuadraPol PSF’s limited FOV extension, unlike these other techniques. Nonetheless, it is important to highlight that the QuadraPol PSF uniquely achieves the elimination of depth ambiguity and the reconstruction without relying on sparsity constraints, distinguishing it from other SVF techniques. Additionally, the QuadraPol PSF provides a better SNR, despite the photon loss due to the linear polarizers (Supplementary Figure S6 and Section 3).
@@ 3.2 Experimental validation using fluorescent beads
We next validated the QuadraPol PSF with fluorescent beads placed on a tilted coverslip (Figure 5(a)). The 3D reconstruction is shown in Figure 5(b), where the reconstruction algorithm accurately resolves the fluorescent beads on the coverslip without noticeable inaccuracies. Additionally, we quantified the full width at half maximum (FWHM) values for the reconstructed images of 72 fluorescent beads from a $z$ scan (Figure 5(c)). The FWHM values are $4.7\pm 1.0$ µm and $4.3\pm 0.9$ µm in the $x$ and $y$ directions, respectively, for in-focus beads. The average FWHM remains within 5.0 µm for a $|z|\leq 0.6$ mm. When beads are defocused by 1 mm, the FWHM increases to an average of 6.0 µm. Note that although the resolution metrics are not directly comparable due to the differences between the FWHM and Rayleigh criteria, the trend observed as a function of $z$ is consistent with our simulations (Figure 4(a)). In the $z$ direction, the average FWHM is consistently below 130.4 µm across a DOF of 1.2 mm and increases to an average of 155.7 µm when defocused by 1 mm. We do not observe resolution degradation near the focal plane, unlike what was seen in simulations using theoretical PSFs (Figure 4(b)). This reaffirms that the system’s aberrations actually improve the depth resolution of the QuadraPol PSF.
@@ 3.3 All-in-focus imaging of bacterial colony on sand surfaces
To address this challenge, we use 3D imaging with the QuadraPol PSF to visualize E. coli tagged with mScarlet-I on the sand surface (Figure 6(a)). Multiple FOVs are stitched together using the Microscopy Image Stitching Tool plugin in imageJ [53]. To determine the correct focusing height map (Figure 6(b)) and generate an all-in-focus image (Figure 6(c)) [54, 55], we use a $100\times 100$ pixel window (side length of 0.35 mm) sliding across the entire FOV. The sharpest focus $z$ position for each window is selected based on image contrast, defined as the difference between the 99th and 1st percentile values in each reconstructed $z$-slice (Supplementary Figure S7). The zoomed regions (Figure 6(d)) show that we successfully resolve sharp reconstructions of the bacterial colonies on the sand particles within a $z$ range of 2 mm from the blurred raw images.
@@ 3.4 Three-dimensional imaging of plant roots with neural fields
We performed the reconstruction using the RL deconvolution with both the experimental and retrieved PSFs, as well as our proposed method based on neural fields in Section 2.3. The image volumes with color-coded depth are presented in Figure 7(a). Observations from full FOV images (Supplementary Figure S8) highlighted that for thicker segments of the roots, the deconvolution results with the experimental PSF appear noisier compared to the method using neural fields, particularly for defocus values of $|z|>2$ mm. Meanwhile, deconvolution with the retrieved PSF shows a broader spread in $z$ (Supplementary Figure S9), due to minor mismatches between the experimental and retrieved PSFs that reduce $z$-axis accuracy and precision. The neural field method addresses these issues by maintaining the accuracy and precision of the deconvolution with the experimental PSF while avoiding its SNR problems.
@@
Figure 7: Volumetric imaging of wheat roots using the QuadraPol PSF. (a) Reconstruction using neural fields. Inset shows a photograph captured using a smartphone camera. (b) The raw polarized fluorescence image. (c) The $xy$ view, and (d,e) the $xy$ and $xz$ views of zoomed regions in (a), reconstructed using (i) deconvolution with experimental PSF, (ii) deconvolution with retrieved PSF, and (iii) neural fields. Scale bar: 5 mm in (a,b), 0.5 mm in the zoomed region of (b), and 1 mm in (c-e). Color bar: depth in mm.
@@ 4 Discussion and conclusion
Our proposed method combines the QuadraPol PSF with neural fields to achieve SVF imaging. A major innovation in hardware is the usage of the four-polarization custom polarizer, which simultaneously encodes depth information across four images (three of which are independent) captured by a polarization camera. This setup maximizes the capabilities of the polarization camera, completely eliminating the estimation ambiguity while maintaining a compact footprint, which allows the QuadraPol PSF to enable SVF without a sparsity constraint and thus achieve a better SNR compared to the state-of-the-art SVF technique. With a single FOV of 3.8 mm $\times$ 4.5 mm, we experimentally demonstrate the ability to capture a $\sim$100 mm3 cubic volume with lateral and axial resolutions of approximately 7 µm and 240 µm, respectively. The extended DOF offered by this system reduces acquisition time by approximately 20 times compared to traditional $z$-scan methods for capturing the same volume. The effectiveness of the QuadraPol PSF is shown in our all-in-focus imaging of bacterial colonies on sand surfaces, where it captures sharp images of most of the colonies in a single snapshot.
@@
For the reconstruction algorithm, neural fields not only demonstrate better performance under noisy conditions but also address limitations of classical reconstruction algorithms. By combining both experimental and retrieved PSFs, our method addresses two key challenges: it overcomes the low SNR encountered in deconvolution using experimental PSFs at larger defocus distances, and it mitigates the reduced deconvolution accuracy resulting from imperfect modeling of the imaging system when using only retrieved PSFs. As demonstrated in our plant root imaging experiments, neural fields produce sharper images compared to deconvolution. The neural field approach also offers additional advantages such as creating a continuous representation of the object and significantly decreasing data storage requirements by an order of magnitude. For instance, storing plant root data as discrete images with 81 axial slices requires approximately 62 GB, whereas neural fields of the same data requires only 6.12 GB with the same data precision.
SVF imaging using the QuadraPol PSF and neural fields provides a powerful tool for large-FOV imaging at high spatial-temporal resolutions. The design principle can be easily adapted to meet specific application requirements. For instance, our current setup prioritizes high temporal resolution by using a polarization camera with a large effective pixel size, thus sacrificing spatial resolution. However, systems requiring a higher spatial sampling rate from the detector could use a standard camera combined with polarization optical elements and temporal multiplexing to address this limitation. It can also be integrated with the DOF extension methods used in MiniScope3D [11] and Fourier DiffuserScope [12] to achieve an improved SBP. In summary, our experimental results highlight its potential, particularly for applications such as studying microbial interactions within the rhizosphere. It also paves the way for new developments in hardware and algorithm designs for SVF imaging techniques.
@@
[…]
Supplemental document See Supplemental Document for supporting content.

### v3

@@ 3.1 Simulation evaluation of system performance
We first assessed the resolution of the QuadraPol PSF by simulating images of two closely positioned point sources. To determine the resolution, we applyed the Rayleigh criterion, which requires at least a 20% dip between the peaks of the two reconstructed spots. We used a high spatial sampling rate to evaluate the resolution more accurately. In the lateral direction, the QuadraPol PSF achieves an in-focus resolution of 4.3 µm. However, this resolution rapidly degrades to approximately 7 µm with a ±$\pm 0.3$ mm defocus. The resolution remains under 9 µm with a defocus extending up to 1.8 mm (Figure 4(a)). To optimally adapt to these resolutions in our experimental setup, we used a camera sampling rate of 3.7 µm in the object space (with a magnification of 1.875) that corresponds to a resolution limit of 7.4 µm, achieving a DOF of $\sim$4 mm with relatively uniform lateral resolution. This extended DOF is notably longer than the 0.26 mm DOF achievable with a standard PSF, which is due to the smaller effective NA of the QuadraPol PSF.
@@
Additionally, we compared our neural fields reconstruction algorithm with the RL deconvolution. While both algorithms perform similarly under ideal conditions, neural fields exhibit significantly improved robustness with noisy images (Figure 4(a) and Supplementary Figure S2). For instance, at a low signal level where a total of $3.16\times 10^{4}$ photons are captured from both emitters, the RL deconvolution results in a highly noisy reconstruction, whereas the neural fields can still distinctly resolve the two spots and offer a two-fold resolution improvement. This improvement is attributed to the fact that the neural fields tend to favor smooth reconstructions [34, 51, 38, 40], effectively leading to a suppression of noise on the reconstruction process.
@@
The axial resolution within the 4 mm DOF ranges from 230 to 250 µm, except for positions close to the focal plane ($|z|<0.1$ mm), where it degrades to 405 µm (Figure 4(b)). This degradation is attributed to the symmetry which causes gradual depth gradient of the in-focus QuadraPol PSF; both $+z$ and $-z$ defocus result in an expansion of the PSF. However, this symmetry is disrupted in the presence of aberrations; when aberrated PSFs were simulated, an improvement in axial resolution near the focal plane was observed (Supplementary Figure S3). We also experimentally demonstrated this improvement in Section 3.2. Similar to the case of lateral direction, using neural fields leads to a better axial resolution compared to RL deconvolution under noisy conditions, although the improvement is less pronounced than in the lateral direction, particularly at very low signal levels. This is likely due to the design of the network’s feature space, which provides more degrees of freedom in the lateral direction than in the axial direction, given the lower axial resolution of the imaging system compared to its lateral resolution. While such a configuration can significantly accelerate the algorithm and reduce memory usage, it can also limit the improvements in axial resolution. Again, we emphasize that this comparison is under ideal conditions and does not account for forward model mismatches that may occur in experimental settings. The space-bandwidth product (SBP) of our system is $\sim$5.2 million voxels over the 3.8 mm $\times$ 4.5 mm FOV and 4 mm depth range. Note that the SBP for the biological experiments in later sections is greater than the value reported here. We are not confined to the 4 mm depth range, as we can still reconstruct the object with slightly reduced resolution at greater defocus distances.
@@
One key advantage of the QuadraPol PSF is its ability to resolve denser objects more effectively. By separating polarization channels, the compact QuadraPol PSF reduces the mixing of information from different 3D positions, thus relaxing the sparsity constraint in the reconstruction algorithm compared to other SVF techniques such as the Miniscope3D [13]. To demonstrate this, we simulated three different line structures and generated images using both the Miniscope3D PSF and the QuadraPol PSF, analyzing them with the same RL deconvolution algorithm without sparsity constraints. For a small structure where the raw image using the Miniscope3D shows minimal overlap, both methods accurately resolve the structure (Figure 4(c)). Additionally, the reconstruction quality of the Miniscope3D for defocused axial positions is superior due to its extended DOF. However, for larger objects where there is substantial information mixture in the raw Miniscope3D images (Figure 4(d,e)), it struggles to accurately recover the structure. In contrast, the QuadraPol PSF still resolves the object effectively with significantly fewer artifacts. This enhanced performance is similarly observed with other dense objects (Supplementary Note S3).
@@
A comparison of the three aspects mentioned above between the QuadraPol PSF and other SVF techniques is shown in Figure 4(h). The most significant advantage of the QuadraPol PSF, which we would like to emphasize, is its ability to eliminate depth ambiguity and reconstruct 3D scenes without relying on sparsity constraints, distinguishing it from other SVF techniques such as MiniScope3D, Fourier DiffuserScope, and PS2F. Additionally, compared to the Fourier light field microscope, which shares the aforementioned advantages, the QuadraPol PSF achieves a higher SBP as it does not compromise the field of view (FOV). Furthermore, the QuadraPol PSF provides a better SNR, despite the photon loss due to the linear polarizers (Supplementary Note S3).
@@ 3.2 Experimental validation using fluorescent beads
We next validated the QuadraPol PSF with fluorescent beads placed on a tilted coverslip (Figure 5(a)). The 3D reconstruction is shown in Figure 5(b), where the reconstruction algorithm accurately resolves the fluorescent beads on the coverslip without noticeable inaccuracies. Quantitative analysis of depth accuracy is provided in Supplementary Figure S5. Additionally, we quantified the full width at half maximum (FWHM) values for the reconstructed images of 72 fluorescent beads from a $z$ scan (Figure 5(c)). The FWHM values are $4.7\pm 1.0$ µm and $4.3\pm 0.9$ µm in the $x$ and $y$ directions, respectively, for in-focus beads. The average FWHM remains within 5.0 µm for a $|z|\leq 0.6$ mm. When beads are defocused by 1 mm, the FWHM increases to an average of 6.0 µm. Note that although the FWHM is not directly comparable to the resolution determined using Rayleigh criteria in Section 3.1, the trend observed as a function of $z$ is consistent with our simulations (Figure 4(a)). In the $z$ direction, the average FWHM is consistently below 130.4 µm across a DOF of 1.2 mm and increases to an average of 155.7 µm when defocused by 1 mm. We do not observe resolution degradation near the focal plane, unlike what was seen in simulations using theoretical PSFs (Figure 4(b)). This reaffirms that the system’s aberrations actually improve the depth resolution of the QuadraPol PSF. Additional resolution analysis (Supplementary Figure S6) based on the Rayleigh criterion using synthesized images from bead data is consistent with the simulation results shown in Figures 4(a,b).
@@ 3.3 All-in-focus imaging of bacterial colony on sand surfaces
To address this challenge, we use 3D imaging with the QuadraPol PSF to visualize E. coli tagged with mScarlet-I on the sand surface (Figure 6(a)). Multiple FOVs are stitched together using the Microscopy Image Stitching Tool plugin in imageJ [56]. To determine the correct focusing height map (Figure 6(b)) and generate an all-in-focus image (Figure 6(c)) [57, 58], we use a $100\times 100$ pixel window (side length of 0.35 mm) sliding across the entire FOV. The sharpest focus $z$ position for each window is selected based on image contrast, defined as the difference between the 99th and 1st percentile values in each slice in the reconstructed $z$-stack (Supplementary Figure S7). The zoomed regions (Figure 6(d)) show that we successfully resolve sharp reconstructions of the bacterial colonies on the sand particles from blurred raw images across the entire sample area, which has a depth variation of 2 mm.
@@ 3.4 Three-dimensional imaging of plant roots with neural fields
We performed the reconstruction using the RL deconvolution with both the experimental and retrieved PSFs, as well as our proposed method based on neural fields in Section 2.3. For the neural fields method, we used experimental PSF for $|z|\leq 2$ mm and retrieved PSF otherwise during the two-step optimization process (Supplementary Note S4). This is due to the degraded quality of experimental PSFs from the low SNR at large defocus distances. In contrast, using both PSFs in a single deconvolution results in sharp discontinuities at the $z$ position transitioning between experimental and retrieved PSFs (Supplementary Note S4). The image volumes with color-coded depth are presented in Figure 7(a). Observations from full FOV images (Supplementary Figure S8) highlighted that for thicker segments of the roots, the deconvolution results with the experimental PSF appear noisier compared to the method using neural fields. Meanwhile, deconvolution with the retrieved PSF shows a broader spread in $z$ (Supplementary Figure S9), due to minor mismatches between the experimental and retrieved PSFs that reduce $z$-axis accuracy and precision. The neural field method addresses these issues by maintaining the accuracy and precision of the deconvolution with the experimental PSF while avoiding its SNR problems.
@@
Figure 7: Volumetric imaging of wheat roots using the QuadraPol PSF. (a) Reconstruction using neural fields. Inset shows a photograph captured using a smartphone camera. (b) The raw polarized fluorescence image. (c) The $xy$ view, and (d,e) the $xy$ and $xz$ views of zoomed regions in (a), reconstructed using (i) deconvolution with experimental PSF, (ii) deconvolution with retrieved PSF, and (iii) neural fields. Scale bar: 5 mm in (a,b), 0.5 mm in the zoomed region of (b), and 1 mm in (c-e). Color bar: depth in mm. Note that the color in the $xy$-view is composed of the entire reconstructed volume, reflecting the intensity contribution from all $z$ slices; this view predominantly displays a color corresponding to the central height of the root.
@@ 4 Discussion and conclusion
Our proposed method combines the QuadraPol PSF with neural fields to achieve SVF imaging. A major innovation in hardware is the usage of the four-polarization custom polarizer, which simultaneously encodes depth information across four images (three of which are independent) captured by a polarization camera. This setup maximizes the capabilities of the polarization camera, completely eliminating the estimation ambiguity while maintaining a compact footprint, which allows the QuadraPol PSF to enable SVF without a sparsity constraint and thus achieve a better SNR compared to the current state-of-the-art SVF techniques. With a single FOV of 3.8 mm $\times$ 4.5 mm, we experimentally demonstrate the ability to capture a $\sim$100 mm3 cubic volume with lateral and axial resolutions of approximately 7 µm and 240 µm, respectively. The extended DOF offered by this system reduces acquisition time by approximately 20 times compared to traditional $z$-scan methods for capturing the same volume. The effectiveness of the QuadraPol PSF is shown in our all-in-focus imaging of bacterial colonies on sand surfaces, where it captures sharp images of most of the colonies in a single snapshot.
@@
For the reconstruction algorithm, neural fields demonstrate better performance compared to deconvolution, as demonstrated in our plant root imaging experiments where neural fields produce sharper images. The neural field approach also offers additional advantages such as creating a continuous representation of the object and significantly decreasing data storage requirements by an order of magnitude. For instance, storing plant root data as discrete images with 81 axial slices requires approximately 62 GB, whereas neural fields of the same data requires only 6.12 GB with the same data precision. Furthermore, this physics-based method is free from any pertaining or any dataset collection. It can also be easily adapted to different system settings.
SVF imaging using the QuadraPol PSF and neural fields provides a powerful tool for large-FOV imaging at high spatial-temporal resolutions. The design principle can be easily adapted to meet specific application requirements. For instance, our current setup prioritizes high temporal resolution by using a polarization camera with a large effective pixel size, thus sacrificing spatial resolution. However, systems requiring a higher spatial sampling rate from the detector could use a standard camera combined with polarization optical elements and temporal multiplexing to address this limitation. One current limitation of the QuadraPol PSF is that it relies on the expansion of four simultaneously-in-focus spots to estimate depth, which inherently limits the imaging depth. If the source-to-focal-plane distance increases, the PSFs in all four polarization channels expand, causing photons to spread across a larger area and consequently reducing the SNR at large defocus distances. In other words, due to the limited DOF extension of the QuadraPol PSF, its SBP still has room for improvement compared with the MiniScope3D [13] or the Fourier DiffuserScope [14] (Figure 4(h)). Integrating the QuadraPol PSF with DOF extension methods used in these other systems, such as using an engineered metasurface to create PSFs in four different polarization channels with distinct focal planes, should improve the SBP. Further, these methods can potentially overcome the limitation of low depth resolution near the focal plane without introducing aberrations. We anticipate that these sophisticated improvements of QuadraPol would be worth implementing as this technology matures. Given the presence of a polarization camera in the QuadraPol PSF, a natural extension of the method would be to simultaneously measure polarization [59]. Similar to what has been demonstrated with the MVR microscope [45], the QuadraPol PSF should exhibit high polarization sensitivity. Previous work in polarization imaging, such as 3D Mueller matrix imaging [60, 61, 62] and Stokes-correlometry [63], has established frameworks for quantitative analysis of light polarization. This direction holds particular promise, as fluorescence polarization has been shown to be critical in plant studies [64, 65] and other areas of biological research [66].
In summary, our experimental results highlight the potential of SVF imaging using the Quadra-Pol PSF and neural fields, particularly for applications such as studying microbial interactions within the rhizosphere. Additionally, we expect that our SVF system can be applied across various fields, including clinical and biomedical research. For instance, we tested our method on synthetic lymph node vasculature data [67] (Supplementary Figure S12), experimental mouse kidney data (Supplementary Figure S13), and demonstration of 3D particle image velocimetry (Supplementary Figure S14 and Supplementary Video S1). The simplicity of the optical setup also holds the potential for miniaturization, which would be useful for in vivo imaging in freely behaving animals. Moreover, in terms of technological advancement, our system introduces a novel approach to SVF design: instead of mapping 3D objects to 2D images, our method maps 3D objects to 3D measurements, with polarization serving as the third dimension. We anticipate this innovative approach will inspire future developments in SVF imaging system designs.